OpenStudy (anonymous):

Find a pair of factors for each number by using the difference between two squares. a. 45 b. 77 c. 112

jimthompson5910 (jim_thompson5910):

a) Notice how 45 = 9*5 45 = (7+2)*(7-2) 45 = 7^2 - 2^2 45 = 49 - 4 ------------- So going backwards, you can say 45 = 49 - 4 45 = 7^2 - 2^2 45 = (7+2)*(7-2) 45 = 9*5

OpenStudy (anonymous):

ok so what do i do next

jimthompson5910 (jim_thompson5910):

Use this technique to do parts b) and c) Let me know what you get

OpenStudy (anonymous):

hold on i will be right back

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

does it make a difference if you go backwards or not also

jimthompson5910 (jim_thompson5910):

I'm just showing you how I got the stuff under the "------" line. You don't have to go backwards if you know how to go forwards (if that makes any sense)

OpenStudy (anonymous):

oh ok yes it does ok hold on

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

could you help

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

still on 77... :(

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

77 = 11*7 77 = (9+2)*(9-2) 77 = 9^2 - 2^2 77 = 81 - 4 ------------------ Now go backwards 77 = 81 - 4 77 = 9^2 - 2^2 77 = (9+2)*(9-2) 77 = 11*7 So the idea is that you work out the first part on scratch paper. Then the second part is what you do on your hw

OpenStudy (anonymous):

ok i am stuck on the (+) (-) part on 112

OpenStudy (anonymous):

on the first part so far i have 112= 8*14

jimthompson5910 (jim_thompson5910):

so 112 = 8*14 112 = (11-3)*(11+3)

OpenStudy (anonymous):

ok so i got 112=8*14 112=(11+3)(11-3) 112=11^3- 3^3 112=116-4

jimthompson5910 (jim_thompson5910):

you square, not cube

jimthompson5910 (jim_thompson5910):

so your 3rd step should be 112 = 11^2 - 3^2

OpenStudy (anonymous):

oh ok soo what do i do next

jimthompson5910 (jim_thompson5910):

11^2 = 121, 3^2 = 9 so 112 = 11^2 - 3^2 becomes 112 = 121 - 9

jimthompson5910 (jim_thompson5910):

Now go backwards: start with 112 = 121 - 9 and work your way up to find the factorization for 112

OpenStudy (anonymous):

112=121-9 112=11^2-3^2 112=(11-3)*(11+3) 112=8*14

jimthompson5910 (jim_thompson5910):

good, you nailed it

OpenStudy (anonymous):

ok so what do i put as the answer? the whole things we just worked on

jimthompson5910 (jim_thompson5910):

what you just wrote is your answer for part c)

jimthompson5910 (jim_thompson5910):

the other parts are similar (just use different numbers)

OpenStudy (anonymous):

so a. is all that with 45 and b is all that with 77 and c. is all that with 112 is the answers?

jimthompson5910 (jim_thompson5910):

yes pretty much, half is scratch work so you know how to start each one

jimthompson5910 (jim_thompson5910):

but you might as well include it

OpenStudy (anonymous):

alrighty Thanks so much

jimthompson5910 (jim_thompson5910):

yw